OpenStudy (empty):

Evaluate the derivative:

OpenStudy (empty):

\[\frac{d}{dn} \left( \sum_{k=0}^n x^k \right)\]

OpenStudy (anonymous):

now the thing is should i consider x as constant or variable of n ?

OpenStudy (anonymous):

not used to x with n xD

OpenStudy (anonymous):

\( \Large \frac{d}{dn} \left( \sum_{k=0}^n x^k \right) \\ \Large \sum_{k=0}^n x^k =\dfrac{1+x^n}{1-x}\\ \Large \dfrac{d}{dn} \left( \dfrac{1+x^n}{1-x} \right) \)

OpenStudy (anonymous):

xD

OpenStudy (empty):

We could do something like this: \[\frac{S(n+1)-S(n)}{(n+1)-n} = \frac{\sum_{k=0}^{n+1} x^k - \sum_{k=0}^n x^k}{1} = x^{n+1}\]

OpenStudy (anonymous):

oh this makes more sense

OpenStudy (irishboy123):

.